# On the effects of radiation on mass transfer in binary stars

### v1

@@ 6 Summary and discussion
  • If the donor envelope is dominated by gas pressure, $P_{\rm gas}\gg P_{\rm rad}$ or equivalently $\Gamma_{\rm Edd}\ll 1$, our model predicts a MT rate comparable to standard models by Jackson et al. (2017) and Kolb & Ritter (1990) in the optically thin and thick regimes, respectively. This is illustrated for a solar-like donor on the red giant branch in Fig. 8.
  • If the donor envelope is dominated by radiation pressure, $P_{\rm rad}\gg P_{\rm gas}$ and $\Gamma_{\rm Edd}\lesssim 1$, our optically thick rates are comparable to predictions of Kolb & Ritter (1990) and later works based on similar assumptions (Fig. 5). The reason is that radiation pressure is already accounted for in the hydrostatic stellar structure used as input in the existing models. However, for Roche-lobe-underfilling donors, our optically thin MT rates are significantly higher than the predictions of previous models of Ritter (1988) or Jackson et al. (2017) (Fig. 5). The increase in the MT rate is exponential in $1-\Gamma_{\text{Edd}}$ and the effect can be included by a simple modification of existing prescriptions (equation 36).
@@
  • Super-Eddington convectively inefficient subsurface layers in single non-rotating donors lead to envelope inflation and density inversions, but should not develop super-Eddington outflows, because the implied mass-loss rates exceed the photon tiring limit (Owocki et al., 2004). In a binary system, the L1 point effectively brings the gravitational potential infinity much closer to the super-Eddington layer and effectively increases the photon tiring limit (equation 48). As a result, a super-Eddington-boosted MT could occur (Fig. 6). By postprocessing a binary evolutionary trajectory, we proposed that super-Eddington-boosted MT can occur before the donor overfills its Roche lobe and that MT rates reach $\gtrsim 10^{-2}\,\rm{M}_{\odot}\,\text{yr}^{-1}$ (Fig. 7). Consequently, this new mode of MT can fundamentally alter the evolutionary trajectories of massive binaries.
@@ 6.1 Implications
We have identified two new effects of radiation pressure on MT: the exponential increase in the optically thin MT rate for donors with a high but sub-Eddington surface $\Gamma_{\text{Edd}}$ and the circumvention of the photon tiring limit in donors possessing super-Eddington convectively inefficient subsurface layers. The importance of the increased optically thin MT rate for binary evolution is likely limited, because although the rates are much higher than traditionally assumed, they are still significantly lower than MT rates for Roche-lobe-overflowing donors or the mass lost from a massive donor by stellar winds. However, the enhanced optically thin rate could be important for donors undergoing wind Roche-lobe overflow and for the stability of massive contact binaries, which are already close to overflowing their outer Lagrangian points (L2 and L3). If substantial mass is lost through L2/L3 via an optically thin outflow analogous to a similar process near L1, it would carry away angular momentum, causing the binary to shrink. If this effect is strong enough, it could destabilize the contact binary, potentially leading to a runaway merger with consequences similar to what was observed in a low-mass merger V1309 Sco (e.g., Tylenda et al., 2011; Pejcha, 2014; Pejcha et al., 2017). To assess this impact, future simulations should incorporate our MT model at the L2/L3 points.
@@
The total amount of mass transferred during the super-Eddington boost depends on how the donor structure responds to this type of mass loss. We can imagine two scenarios. Either the super-Eddington MT simply adds to the traditional MT rate. The effective outcome of this could be simply that the Roche-lobe overflow occurs earlier than traditionally predicted, but the ultimate fate of the binary remains unaffected. Alternatively, the high $\dot{M}_{\text{out,L1}}$ induces a strong response in the donor and the super-Eddington convectively inefficient layer disappears and the donor shrinks. The super-Eddington layer could reconstitute again, leading to expansion and repeating of the super-Eddington-boosted MT. In such a case, the donor might completely avoid overfilling its Roche lobe, and the binary evolution would be significantly altered. However, it is not clear how much different the ultimate fate of the binary would be.
@@
This last scenario motivates further speculation on the cause of LBV outbursts. The connection between these eruptions and super-Eddington outflows has a long history (e.g., Smith & Owocki, 2006), but it has not been clear what triggers the enhanced super-Eddington state. The scenario in Fig. 7 might provide a clue to answer this question. When a subsurface super-Eddington layer is unable to drive an outflow due to photon tiring, it becomes subject to instabilities leading to up and down flows as well as small eruptions (e.g., Owocki et al., 2004; Jiang et al., 2015, 2018). Once the donor slowly expands sufficiently close to the L1 point, the photon tiring limit is relaxed and the donor is able to drive an outflow through the L1 nozzle. Envelope instabilities are likely to make this process manifest as a series of distinct irregular bursts rather than a smooth outflow. The expected mass-transfer/outflow rates are consistent with what was inferred for LBV outbursts (Smith, 2014). Depending on how the donor responds to this type of mass loss, there might be multiple phases of activity separated by inactive periods during which the driving region in the donor reconstitutes. In general, the donor might lose up to several $\rm{M}_{\odot}$ of material before the conditions for the outflow are no longer satisfied.
@@
The diversity of outcomes of this scenario for LBV outbursts can be further increased by considering super-Eddington layers due to hydrogen, helium, or iron ionization, accretion onto the companion, which might be a compact object able to accelerate part of the material to high velocities, or eccentric passages of the companion, where the effective L1 point passes close to the surface only temporarily. All of these issues are nontrivially mapped on the properties of the initial conditions of binary components and their orbit. However, this scenario cannot possibly explain all LBV outbursts. For example, since the driving mechanism ultimate relies on stellar photons, the kinetic energy of the ejected material is limited by the energy radiated by the star, which is violated in the brightest LBV eruptions such as $\eta$ Car (Smith et al., 2011).
@@ 6.2 Limitations and future work
We now discuss several limitations of our work and how we can improve our results in the future. The key assumption in our model is the von Zeipel theorem. Although this theorem is valid in stellar interiors and for slowly rotating stars, it becomes less reliable near the stellar surface and Lagrangian points. Moreover, it is inherently inconsistent. In a stationary system where the radiative flux governs energy transport, energy conservation dictates a divergence-free flux (equations 1d, 1e). However, the assumption of proportionality between radiative flux and effective gravity (equation 16) implies the existence of the Poisson equation in the donor envelope, $\Delta\phi_{\rm R}=0$, in regions where $L(\phi_{\rm R})\approx\rm{const.}$ and $M(\phi_{\rm R})\approx\rm{const}$. However, in reality, $\Delta\phi_{\rm R}=-2\omega^{2}$, where $\omega$ is the binary orbital frequency. This inconsistency was resolved in 2D by Espinosa Lara & Rieutord (2011), which suggests a promising avenue for adopting their approach along the binary axis and for further generalization to 3D.
@@
Our model for the super-Eddington MT boost in Section 5.1 assumes that the material will flow around L1 once the photon tiring limit becomes relaxed there. However, this is only a necessary but not sufficient condition. Instead, it is possible that lateral structuring or “porosity” will reduce the effective coupling between radiation and matter similar to what was previously argued for spherical stars (e.g., Shaviv, 2001a; Owocki et al., 2004), leading to a time-dependent but still nearly hydrostatic structure (Jiang et al., 2015, 2018).
@@
There are several ways to further improve our model and understanding of MT in binaries in general. As a first step, 1D binary evolution models that include modifications to the optically thin MT and some prescription for the possible super-Eddington boost would be useful to investigate self-consistent response of the binary to these new modes of MT and to assess the effect on the outcomes of binary evolution and to connect to phenomena like LBV outbursts. However, the model of super-Eddington boost presented in Section 5 depends on several multidimensional radiation-hydrodynamic phenomena: von Zeipel theorem in Roche geometry, photon porosity of radiation pressure-dominated layers, inefficient convection, and the Coriolis force. Although effective 1D theories exist for each of these phenomena separately, capturing the mutual non-linear interactions in the context of some new 1D theory of super-Eddington MT boost is unlikely to be successful. Instead, future multidimensional, time-dependent radiation hydrodynamic simulations of MT through L1 are essential to test the validity of the simple model presented here and to further clarify the flow geometry, MT rate, and response to the super-Eddington condition (e.g., Jiang et al., 2015, 2018; Dickson, 2024). Given the difficulty of capturing 3D phenomena with 1D models, the relatively small but noticeable differences in the existing 1D MT prescriptions (Fig. 4) suggest that using a simple analytic prescription such as the one presented here in Section 4.2 might do as good a job as some more complicated 1D models.
@@ Acknowledgements
JC thanks Damien Gagnier for useful comments. This research was supported by Czech Science Foundation Grant No. 24-11023S.
@@ Appendix A Derivation of 3D velocity equation
Thus, using the definitions of $q_{\rm rad}$ and $\boldsymbol{f}_{\rm rad}$ in equations (7,8) gives the 3D velocity equation for a general EOS that reads (cf. Lubow & Shu 1975, eq. 10)
@@ Appendix B Derivation of 1D radiation hydrodynamic equations and 1D velocity equation
$-\frac{c\lambda}{\rho\kappa_{\rm R}}\frac{{\rm d}E_{\rm rad}}{{\rm d}x}=\frac{L_{\rm N}}{Q}-\left(\rho\epsilon^{*}+P_{\rm gas}+\frac{3-f}{2}E_{\rm rad}\right)v.$ (61)
@@ Appendix C Our previous approach and other models
Here, we show the 1D hydrodynamic equations governing the gas flow in our original model (CP23) for convenience. The equations read
@@
The optically thin MT rate of Jackson et al. (2017) is given by
$-\dot{M}_{\rm d}\equiv\dot{M}_{\rm J}=\dot{M}_{\rm{J,0}}\exp\left(-\frac{\phi_{1}-\phi_{\rm ph}}{c_{T,\rm gas}^{2}}\right),\quad\dot{M}_{\rm{J,0}}=\frac{2\pi}{\sqrt{\rm e}}\frac{1}{\sqrt{BC}}c_{T,\rm gas}^{3}\rho_{\rm ph},$ (66)
@@
where the volume-equivalent radius $r_{V}$ is defined using a sphere with a volume $V$ enclosed by the equipotential surface on the donor’s side of the L1 plane, $V(\phi_{V})=4\pi r_{V}^{3}/3$. The optically thick MT rate of Kolb & Ritter (1990) is given by
$-\dot{M}_{\rm d}\equiv\dot{M}_{\rm{KR}}=\dot{M}_{\rm{J,0}}+\frac{2\pi}{\sqrt{BC}}\int_{P_{\rm ph}}^{\bar{P}(R_{\rm L})}F_{3}\left(\gamma\right)c_{T,\rm gas}{\rm d}\bar{P},\quad\text{where}\quad F_{3}\left(\gamma\right)=\gamma^{\frac{1}{2}}\left(\frac{2}{\gamma+1}\right)^{\frac{\gamma+1}{2\left(\gamma-1\right)}},$ (68)
@@ Appendix D $1\>\rm{M}_{\odot}$ donor on the red giant branch
In Fig. 8, we compare $\dot{M}_{\rm an,thin}(\Delta R_{\rm d})$ and $\dot{M}_{\rm an,thick}(\Delta R_{\rm d})$ with the models of Jackson et al. (2017) ($\dot{M}_{\rm J}$) and Kolb & Ritter (1990) ($\dot{M}_{\rm KR}$). The photospheric pressure scale height in this case is $H_{P,\rm ph}=0.011R_{\rm d}=0.96\,\rm{R}_{\odot}$. We see, as expected, that for a gas pressure dominated envelope, $P_{\rm gas}\gg P_{\rm rad}$, the differences between the two models are small (to within 20 %).
Figure 8: Comparison of our analytical MT prescription $\dot{M}_{\rm an,thin}$ and $\dot{M}_{\rm an,thick}$ with the optically thin prescription introduced by Jackson et al. (2017), $\dot{M}_{\rm J},$ and the optically thick prescription introduced by Kolb & Ritter (1990), $\dot{M}_{\rm KR}$. We also show the total Eddington factor $\Gamma_{\rm tot}=L/L_{\text{Edd}}$ and the Eddington factor $\Gamma_{\rm Edd}$ (equation 28) for comparison. The case shown is for an isolated $1\rm{M}_{\odot}$ solar-like star on the red giant branch.

### v2

@@ 6 Summary and discussion
  • If the donor envelope is dominated by gas pressure, $P_{\rm gas}\gg P_{\rm rad}$ or equivalently $\Gamma_{\rm Edd}\ll 1$, our model predicts a MT rate comparable to standard models by jackson2017 and kolb1990 in the optically thin and thick regimes, respectively. This is illustrated for a solar-like donor on the red giant branch in Fig. 8.
  • If the donor envelope is dominated by radiation pressure, $P_{\rm rad}\gg P_{\rm gas}$ and $\Gamma_{\rm Edd}\lesssim 1$, our optically thick rates are comparable to predictions of kolb1990 and later works based on similar assumptions (Fig. 5). The reason is that radiation pressure is already accounted for in the hydrostatic stellar structure used as input in the existing models. However, for Roche-lobe-underfilling donors, our optically thin MT rates are significantly higher than the predictions of previous models of ritter1988 or jackson2017 (Fig. 5). The increase in the MT rate is exponential in $1-\Gamma_{\text{Edd}}$ and the effect can be included by a simple modification of existing prescriptions (equation 36).
@@
  • Super-Eddington convectively inefficient subsurface layers in single non-rotating donors lead to envelope inflation and density inversions, but should not develop super-Eddington outflows, because the implied mass-loss rates exceed the photon tiring limit (owocki2004). In a binary system, the L1 point effectively brings the gravitational potential infinity much closer to the super-Eddington layer and effectively increases the photon tiring limit (equation 48). As a result, a super-Eddington-boosted MT could occur (Fig. 6). By postprocessing a binary evolutionary trajectory, we proposed that super-Eddington-boosted MT can occur before the donor overfills its Roche lobe and that MT rates reach $\gtrsim 10^{-2}\,\rm{M}_{\odot}\,\text{yr}^{-1}$ (Fig. 7). Consequently, this new mode of MT can fundamentally alter the evolutionary trajectories of massive binaries.
@@ 6.1 Implications
We have identified two new effects of radiation pressure on MT: the exponential increase in the optically thin MT rate for donors with a high but sub-Eddington surface $\Gamma_{\text{Edd}}$ and the circumvention of the photon tiring limit in donors possessing super-Eddington convectively inefficient subsurface layers. The importance of the increased optically thin MT rate for binary evolution is likely limited, because although the rates are much higher than traditionally assumed, they are still significantly lower than MT rates for Roche-lobe-overflowing donors or the mass lost from a massive donor by stellar winds. Moreover, the details of MT via the L1 point or the precise value of the Roche-lobe radius are not crucial for binary evolution, provided that the MT is stable and the long-term evolution is driven by other factors such as the nuclear or thermal timescale of the donor, as already pointed out by paczynski1971. In such cases, it is more important to consider mass and angular momentum loss from the binary.
Indeed, the enhanced optically thin rate could be important for donors undergoing wind Roche-lobe overflow and for the stability of massive contact binaries, which are already close to overflowing their outer Lagrangian points (L2 and L3). If substantial mass is lost through L2/L3 via an optically thin outflow analogous to a similar process near L1, it would carry away angular momentum, causing the binary to shrink. If this effect is strong enough, it could destabilize the contact binary, potentially leading to a runaway merger with consequences similar to what was observed in a low-mass merger V1309 Sco (e.g., tylenda2011; pejcha2014; pejcha2017). To assess this impact, future simulations should incorporate our MT model at the L2/L3 points.
@@
The total amount of mass transferred during the super-Eddington boost depends on how the donor structure responds to this type of mass loss. We can imagine two scenarios. First, the super-Eddington MT simply adds to the traditional MT rate. The effective outcome of this could be simply that the Roche-lobe overflow occurs earlier than traditionally predicted, but the ultimate fate of the binary remains unaffected. Alternatively, if the mass transfer is already unstable and a hydrostatic structure can no longer be maintained in the outer layers, this additional boost could further accelerate the instability.
Second, the high $\dot{M}_{\text{out,L1}}$ induces a strong response in the donor and the super-Eddington convectively inefficient layer disappears and the donor shrinks. The super-Eddington layer could reconstitute again, leading to expansion and repeating of the super-Eddington-boosted MT. In such a case, the donor might completely avoid overfilling its Roche lobe, and the binary evolution would be significantly altered. However, it is not clear how much different the ultimate fate of the binary would be.
This last scenario motivates further speculation on the cause of LBV outbursts. LBVs are hot, highly luminous massive stars that undergo large, quasi-periodic variations in brightness, radius, and photospheric temperature on timescales of years to decades, often accompanied by episodes of enhanced mass loss (smith2014). The connection between these eruptions and super-Eddington outflows has a long history (e.g., smith_owocki06), but it has not been clear what triggers the enhanced super-Eddington state.
A single-star scenario was recently proposed by grassitelli2021. In addition to an inflated envelope close to the Eddington limit, they required two additional conditions to trigger the bursting activity: first, a temperature range in which decreasing opacities do not lead to an accelerating outflow; and second, a mass-loss rate that increases with decreasing temperature.
@@
Our results in Fig. 7 might lead to a binary scenario for LBV outbursts. In addition to a super-Eddington, convectively inefficient subsurface layer, we also require this layer to be sufficiently close to the L1 point. When it is far from L1 and unable to drive an outflow due to photon tiring, it becomes subject to instabilities leading to up and down flows as well as small eruptions (e.g., owocki2004; jiang2015; jiang2018). Once the donor slowly expands sufficiently close to the L1 point, the photon tiring limit is relaxed and the donor is able to drive an outflow through the L1 nozzle. Envelope instabilities are likely to make this process manifest as a series of distinct irregular bursts rather than a smooth outflow. The expected mass-transfer/outflow rates are consistent with what was inferred for LBV outbursts (smith2014). Depending on how the donor responds to this type of mass loss, there might be multiple phases of activity separated by inactive periods during which the driving region in the donor reconstitutes. In general, the donor might lose up to several $\rm{M}_{\odot}$ of material before the conditions for the outflow are no longer satisfied.
At first sight, such a scenario might seem extremely unlikely to occur. In this paragraph, we argue that this is not necessarily the case. If we evolve the $30\,\mathrm{M}_{\odot}$ low-metallicity donor from Sec. 5.2 as a single star, the super-Eddington and convectively inefficient subsurface layer is present for $R_{\rm d}/R_{\odot}$ between roughly 540 and 910. Assuming that the donor is in a binary, Keplerian orbit, and adopting the simple approximation for the Roche-lobe radius from paczynski1971, $R_{\rm L}/a=0.46[q/(1+q)]^{1/3}$, we obtain the following expression for the binary orbital period: $P_{\rm orb}=20\sqrt{R_{\rm L}^{3}/GM_{\rm d}}=2.3~\mathrm{yr}~R_{\rm L,540}^{3/2}M_{\rm d,30}^{-1/2}$, where $R_{\rm L,540}=R_{\rm L}/540\,\rm{R}_{\odot}$ and $M_{\rm d,30}=M_{\rm d}/30\,\rm{M}_{\odot}$. This expression does not depend on the binary mass ratio. Thus, if the $30\,\mathrm{M}_{\odot}$ low-metallicity donor resides in a binary with an orbital period between roughly 2.3 and 5.1 years, it should overflow its Roche lobe while possessing a super-Eddington, convectively inefficient subsurface layer, and a super-Eddington MT boost could occur. Given that most massive stars are in binaries (e.g. sana2012), these numbers already suggest that such a scenario might not be extremely rare.
@@
The diversity of outcomes of this binary scenario for LBV outbursts can be further increased by considering super-Eddington layers due to hydrogen, helium, or iron ionization, accretion onto the companion, which might be a compact object able to accelerate part of the material to high velocities, or eccentric passages of the companion, where the effective L1 point passes close to the surface only temporarily. All of these issues are nontrivially mapped on the properties of the initial conditions of binary components and their orbit. However, this scenario cannot possibly explain all LBV outbursts. For example, since the driving mechanism ultimate relies on stellar photons, the kinetic energy of the ejected material is limited by the energy radiated by the star, which is violated in the brightest LBV eruptions such as $\eta$ Car (smith2011).
@@ 6.2 Limitations and future work
We now discuss several limitations of our work and how we can improve our results in the future. The key assumption in our model is the von Zeipel theorem. Although this theorem is valid in stellar interiors and for slowly rotating stars, it becomes less reliable near the stellar surface and Lagrangian points. Moreover, it is inherently inconsistent. In a stationary system where the radiative flux governs energy transport, energy conservation dictates a divergence-free flux (equations 1d, 1e). However, the assumption of proportionality between radiative flux and effective gravity (equation 16) implies the existence of the Poisson equation in the donor envelope, $\Delta\phi_{\rm R}=0$, in regions where $L(\phi_{\rm R})\approx\rm{const.}$ and $M(\phi_{\rm R})\approx\rm{const}$. However, in reality, $\Delta\phi_{\rm R}=-2\omega^{2}$, where $\omega$ is the binary orbital frequency. This inconsistency was resolved in 2D by espinosa-lara2011, which suggests a promising avenue for adopting their approach along the binary axis and for further generalization to 3D.
Another simplification of our 1D approach is the neglect of circulation (due to the one-dimensionality of the method) and, more importantly, the neglect of the Coriolis force. Recently, ryu2025 performed 3D hydrodynamical simulations of MT through the L1 point, including the Coriolis force and assuming an ideal-gas EOS. They found that the Coriolis force affects the stream morphology by reducing the stream’s cross-section and shifting its origin to the donor’s trailing side. However, the resulting MT rates were only mildly suppressed relative to analytic predictions, within a factor of 2 at the L1 point. Without the Coriolis force, their results resemble our nozzle picture and agree with analytic prescriptions to within 5 percent. This suggests that including the Coriolis force in our model could mildly reduce the MT rate (by up to a factor of a few), but this effect should remain subdominant compared to, for example, the super-Eddington MT boost discussed in this work.
@@
Our model for the super-Eddington MT boost in Section 5.1 assumes that the material will flow around L1 once the photon tiring limit becomes relaxed there. However, this is only a necessary but not sufficient condition. Instead, it is possible that lateral structuring or “porosity” will reduce the effective coupling between radiation and matter similar to what was previously argued for spherical stars (e.g., shaviv01b; owocki2004), leading to a time-dependent but still nearly hydrostatic structure (jiang2015; jiang2018).
@@
There are several ways to further improve our model and understanding of MT in binaries in general. As a first step, 1D binary evolution models that include modifications to the optically thin MT and some prescription for the possible super-Eddington boost would be useful to investigate self-consistent response of the binary to these new modes of MT and to assess the effect on the outcomes of binary evolution and to connect to phenomena like LBV outbursts. However, the model of super-Eddington boost presented in Section 5 depends on several multidimensional radiation-hydrodynamic phenomena: von Zeipel theorem in Roche geometry, photon porosity of radiation pressure-dominated layers, inefficient convection, and the Coriolis force. Although effective 1D theories exist for each of these phenomena separately, capturing the mutual non-linear interactions in the context of some new 1D theory of super-Eddington MT boost is unlikely to be successful. Instead, future multidimensional, time-dependent radiation hydrodynamic simulations of MT through L1 are essential to test the validity of the simple model presented here and to further clarify the flow geometry, MT rate, and response to the super-Eddington condition (e.g., jiang2015; jiang2018; dickson2024; ryu2025). Given the difficulty of capturing 3D phenomena with 1D models, the relatively small but noticeable differences in the existing 1D MT prescriptions (Fig. 4) suggest that using a simple analytic prescription such as the one presented here in Section 4.2 might do as good a job as some more complicated 1D models.
@@ Acknowledgements
JC thanks Damien Gagnier for useful comments. This research was supported by Czech Science Foundation Grant No. 24-11023S. JC also acknowledges support by the ETAg grant PRG2159 during manuscript revision.
@@ Appendix A Derivation of 3D velocity equation
Thus, using the definitions of $q_{\rm rad}$ and $\boldsymbol{f}_{\rm rad}$ in equations (7,8) gives the 3D velocity equation for a general EOS that reads (cf. lubow1975, eq. 10)
@@ Appendix B Derivation of 1D radiation hydrodynamic equations and 1D velocity equation
$\frac{c\lambda}{\rho\kappa_{\rm R}}\frac{{\rm d}E_{\rm rad}}{{\rm d}x}=\frac{L_{\rm N}}{Q}-\left(\rho\epsilon^{*}+P_{\rm gas}+\frac{3-f}{2}E_{\rm rad}\right)v.$ (61)
@@ Appendix C Our previous approach and other models
Here, we show the 1D hydrodynamic equations governing the gas flow in our original model (cehula2023) for convenience. The equations read
@@
The optically thin MT rate of jackson2017 is given by
$\dot{M}_{\rm d}\equiv\dot{M}_{\rm J}=\dot{M}_{\rm{J,0}}\exp\left(-\frac{\phi_{1}-\phi_{\rm ph}}{c_{T,\rm gas}^{2}}\right),\quad\dot{M}_{\rm{J,0}}=\frac{2\pi}{\sqrt{\rm e}}\frac{1}{\sqrt{BC}}c_{T,\rm gas}^{3}\rho_{\rm ph},$ (66)
@@
where the volume-equivalent radius $r_{V}$ is defined using a sphere with a volume $V$ enclosed by the equipotential surface on the donor’s side of the L1 plane, $V(\phi_{V})=4\pi r_{V}^{3}/3$. The optically thick MT rate of kolb1990 is given by
$\dot{M}_{\rm d}\equiv\dot{M}_{\rm{KR}}=\dot{M}_{\rm{J,0}}+\frac{2\pi}{\sqrt{BC}}\int_{P_{\rm ph}}^{\bar{P}(R_{\rm L})}F_{3}\left(\gamma\right)c_{T,\rm gas}{\rm d}\bar{P},\quad\text{where}\quad F_{3}\left(\gamma\right)=\gamma^{\frac{1}{2}}\left(\frac{2}{\gamma+1}\right)^{\frac{\gamma+1}{2\left(\gamma-1\right)}},$ (68)
@@ Appendix D $1\>\rm{M}_{\odot}$ donor on the red giant branch
In Fig. 8, we compare $\dot{M}_{\rm an,thin}(\Delta R_{\rm d})$ and $\dot{M}_{\rm an,thick}(\Delta R_{\rm d})$ with the models of jackson2017 ($\dot{M}_{\rm J}$) and kolb1990 ($\dot{M}_{\rm KR}$). The photospheric pressure scale height in this case is $H_{P,\rm ph}=0.011R_{\rm d}=0.96\,\rm{R}_{\odot}$. We see, as expected, that for a gas pressure dominated envelope, $P_{\rm gas}\gg P_{\rm rad}$, the differences between the two models are small (to within 20 %).
Figure 8: Comparison of our analytical MT prescription $\dot{M}_{\rm an,thin}$ and $\dot{M}_{\rm an,thick}$ with the optically thin prescription introduced by jackson2017, $\dot{M}_{\rm J},$ and the optically thick prescription introduced by kolb1990, $\dot{M}_{\rm KR}$. We also show the total Eddington factor $\Gamma_{\rm tot}=L/L_{\text{Edd}}$ and the Eddington factor $\Gamma_{\rm Edd}$ (equation 28) for comparison. The case shown is for an isolated $1\rm{M}_{\odot}$ solar-like star on the red giant branch.